2.2g of a compopund of phosphorous and sulphur has1.24g of P in it  Determine its empirical formula

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

percentage of sulpher = 1.24/2.2 *100 =56.4

percentage of P = 100-56.4 =43.6

relative number of molecules  of S = 56.4/32 = 1.76

relative number of molecules  of P = 43.6 /31 = 1.41

intigral ratio of sulpher is =5

intigral ration of P   = 4

so imperical formula  = P4S5

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin